# Anomaly Detection in Complex Dynamical Systems: A Systematic Framework Using Embedding Theory and Physics-Inspired Consistency

### v1

@@ Abstract
Anomaly detection in complex dynamical systems is essential for ensuring reliability, safety, and efficiency in industrial and cyber-physical infrastructures. Predictive maintenance helps prevent costly failures, while cybersecurity monitoring has become critical as digitized systems face growing threats. Many of these systems exhibit oscillatory behaviors and bounded motion, requiring anomaly detection methods that capture structured temporal dependencies while adhering to physical consistency principles. In this work, we propose a system-theoretic approach to anomaly detection, grounded in classical embedding theory and physics-inspired consistency principles. We build upon the Fractal Whitney Embedding Prevalence Theorem, extending traditional embedding techniques to complex system dynamics. Additionally, we introduce state-derivative pairs as an embedding strategy to capture system evolution. To enforce temporal coherence, we develop a Temporal Differential Consistency Autoencoder (TDC-AE), incorporating a TDC-Loss that aligns the approximated derivatives of latent variables with their dynamic representations. We evaluate our method on the C-MAPSS dataset, a benchmark for turbofan aeroengine degradation. TDC-AE outperforms LSTMs and Transformers while achieving a 200x reduction in MAC operations, making it particularly suited for lightweight edge computing. Our findings support the hypothesis that anomalies disrupt stable system dynamics, providing a robust, interpretable signal for anomaly detection.
@@ I Introduction
Anomaly detection in complex physical dynamical systems is a critical research area as industrial and engineered systems become more sophisticated. Identifying deviations from expected behavior is essential for ensuring reliability, safety, and efficiency. This is particularly relevant in predictive maintenance and cybersecurity monitoring. In industrial systems and rotating machinery, early fault detection helps prevent costly failures and downtime [1]. Meanwhile, as critical infrastructures—such as power grids and water distribution systems—become increasingly digitized, the risk of cyber threats has grown significantly [2, 3]. Recent attacks [4, 5] on cyber-physical systems highlight the need for robust monitoring techniques to detect both malicious intrusions and system failures. Ensuring the security and stability of these dynamical systems requires adaptive anomaly detection methods capable of addressing evolving threats and operational challenges.
@@ II-A Time-Series Modeling
Models like LSTMs and Transformers with attention mechanisms have been successfully applied to time-series anomaly detection in fields such as mechanical enginered system, aerospace, and industrial monitoring [6, 7, 8]. While these methods effectively capture long-range dependencies and irregular temporal patterns, they are computationally expensive. For example, common benchmark datasets require time steps ranging from 30 to 500 with hidden dimensions between 20 and 50, illustrating the complexity of modeling system dynamics efficiently [9]. Their high memory requirements often exceed the constraints of typical MCU-level devices, making real-time deployment infeasible For example, a typical LSTM with 500 time steps and 50 hidden units already surpasses the available memory of common low-power devices, limiting its practical applicability. Additionally, the sequential nature of RNNs restricts parallelization, further increasing computational cost [10]. These challenges highlight the need for alternative approaches that balance computational efficiency with robust anomaly detection, aligning with broader goals of sustainability and practical deployability.
@@ II-B Physics-Informed Methods
In many applications, incorporating domain knowledge can help reduce computational demands. Physics-informed neural networks (PINNs) embed physical laws directly into neural networks, enabling them to leverage known system dynamics [11, 12, 13]. However, applying PINNs to complex dynamical physical systems remains challenging. These systems involve numerous interdependent physical principles, and explicitly modeling them within a neural network would require extensive computational resources and domain expertise, making large-scale applications impractical [14, 15].
@@ II-D Proposed Approach
Instead of relying on overly complex models that attempt to directly differentiate between normal and abnormal behavior, we propose a system-theoretic approach inspired by classical embedding theory to study the dimensionality of the system’s latent representation. Our approach introduces physics-inspired consistency principles that approximate the underlying causal mechanisms governing the system dynamics, without explicitly enforcing physical laws. We hypothesize that complex systems in a stable regime exhibit predictable behavior, allowing for well-approximated lower-dimensional embeddings. In contrast, anomalies introduce additional complexity, disrupting these stable relationships. We aim to leverage this property to detect anomalous states of the system when the learned embedding no longer adequately captures the system dynamics.
@@ III Theoretical Foundation & Methods
In this work, we investigate the mathematical foundations of embedding a bounded dynamical system, such as a rotary machine, into a latent space using autoencoders. Given a system governed by well-defined dynamical laws and a set of measurement values recorded as a time series, our objective is to establish conditions under which the system’s dynamics can be reliably mapped to the latent space. By leveraging the learned manifold structure, we aim to detect deviations from the normal operating state within the latent space and subsequently infer corresponding deviations in the measurement domain and, ultimately, in the physical system. This formulation aims to provide a rigorous framework for anomaly detection and system health monitoring based on latent space representations.
III-A General Definition of an Oscillatory Dynamical System
$\dot{\mathbf{x}}=\mathbf{f}(\mathbf{x},t),$ (1)
$\ddot{\theta}+g(\theta,\dot{\theta})+h(\theta,t)=0,$ (2)
@@ III-A General Definition of an Oscillatory Dynamical System
This formulation illustrates that the state variables of an oscillatory system, the angular displacement $\theta$ and its derivative $\dot{\theta}$, define a phase space in which the system’s behavior can be analyzed. In this phase space, it is possible to identify bounded regions corresponding to normal oscillations, where the system remains confined to a predictable, periodic trajectory. When the system leaves this bounded region, it indicates the onset of unbounded oscillations or deviations from normal behavior.
The key question is whether this concept extends to complex dynamical systems and what constraints must be addressed for a successful transition. Specifically, we seek the mathematical and structural conditions for mapping system dynamics to a reduced representation while preserving phase space properties.
High-dimensional, nonlinear systems introduce challenges such as maintaining topological consistency and preserving dynamical coherence. The goal is to identify the fundamental conditions and assumptions governing this embedding.
We systematically examine the mathematical foundation and computational requirements for such mappings, informing the design of effective latent representations in an autoencoder architecture for detecting deviations in high-dimensional systems
III-B Embedding Complex System Dynamics into a Latent Space
We begin by formalizing the mathematical foundation for embedding high-dimensional dynamical systems into a reduced representation. Consider a complex bounded dynamical system with state variables $x\in\mathbb{R}^{\pi}$, where $\pi$ represents the dimension of the physical system and is typically much larger than the number of available measurements $(\pi\gg k$).
@@ III-B Embedding Complex System Dynamics into a Latent Space
In the normal state, system trajectories should be mapped onto a compact manifold in the latent space. The goal is to leverage the embedded representation so that a deviation in the latent space reliably indicates a departure from the stable state at the physical level. We first analyze the latent space encoding of a dynamical system from a system theory perspective. While system dynamics in a stable regime often evolve on a structured manifold, measurements are not necessarily taken from a smooth manifold. These manifolds, in the mathematical sense, locally resemble Euclidean space $\mathbb{R}^{n}$ and allow for the application of calculus [22]. However, real-world measurements may be noisy or lie on fractal-like sets, requiring a more general mathematical framework.
Therefore, we turn to the Fractal Whitney Embedding Prevalence Theorem, which extends the classical Whitney and Takens embedding results to settings where the underlying structure is not a smooth manifold. A key aspect of this generalization is the notion of dimensionality for compact subsets, which leads to the definition of box-counting dimension.
@@
For a positive number $\varepsilon$, let $A_{\varepsilon}$ be the set of all points within $\varepsilon$ of $A$, i.e.,
$A_{\varepsilon}=\{x\in\mathbb{R}^{n}:\|x-a\|\leq\varepsilon\text{ for some }a\in A\}.$
Let $\operatorname{vol}(A_{\varepsilon})$ denote the $n$-dimensional outer volume of $A_{\varepsilon}$. The box-counting dimension of $A$ is then defined as
$\operatorname{boxdim}(A)=n-\lim_{\varepsilon\to 0}\frac{\log\operatorname{vol}(A_{\varepsilon})}{\log\varepsilon}.$
@@
If the measurement function is well-behaved, we can approximate the box-counting dimension using the dimension of bounded physical quantities that govern the real system dynamics in a stable regime. This allows us to estimate the required embedding dimension $n$ to achieve a faithful description of the system’s dynamics.
@@
To effectively utilize this type of embedding, we require more than just a one-to-one mapping—we also need consistent differential information. That is, small changes in the latent space should correspond to meaningful changes in the measurement space. A fundamental tool for analyzing smooth mappings between differentiable manifolds is the Jacobian matrix, which captures the local behavior of the map in terms of its partial derivatives. The immersion theorem ensures that local neighborhoods in the latent space retain the differential structure of the original system. This guarantees that the mapping not only remains locally invertible but also that variations in the system state are meaningfully preserved in the latent representation.
Theorem 2 (Immersion [22]).
Suppose $F:M\to N$ is a smooth map and $p\in M$. If $dF_{p}$ is injective, then there exists a neighborhood $U$ of $p$ such that the restriction $F|_{U}$ is an immersion.
A sufficient condition for the immersion property is that the Jacobian of $F$ has full column rank, ensuring that $dF_{p}$ is injective [22]. By establishing that our encoder is a one-to-one map and an immersion within the considered subspace of the system’s stable regime, we create a meaningful latent representation. From a theoretical perspective, this allows us to associate deviations in the latent space with deviations in the physical system.
III-C Application to Autoencoder-based Representations
Having established the theoretical framework, we now seek to apply it to a practical use case. This requires two key components. First, we need a bounded dynamical system that operates in a stable regime with accessible measurement values. For this, we use the NASA C-MAPSS dataset, which provides time-series sensor data from aircraft engines—a complex yet stable dynamical system suitable for our approach.
Second, we need to develop an algorithm that trains an autoencoder to learn embeddings aligned with our theoretical framework. This involves constructing a latent space representation based on state-derivative pairs while ensuring a one-to-one mapping and an immersion within the system’s normal operating regime.
III-C1 C-MAPSS Dataset
Figure 3: Simplified C-MAPSS engine model with five rotating components: Fan, LPC, HPC, HPT, and LPT. N1 represents the low-pressure spool (Fan, LPC, LPT), while N2 represents the high-pressure spool (HPC, HPT).
The NASA Commercial Modular Aero-Propulsion System Simulation (C-MAPSS) dataset [24] is a widely used benchmark in prognostics and health management studies [25, 26, 27] . It consists of multivariate time series data representing the operational history of a fleet of engines, each experiencing different initial wear and manufacturing variations. The dataset includes three operational settings that significantly affect engine performance, as well as sensor readings capturing engine conditions, which are subject to noise contamination. Each engine operates normally at the beginning of the time series before developing a fault over time.
@@ III-C1 C-MAPSS Dataset
In this study, we specifically use the FD001 subset of the C-MAPSS dataset, which contains data from 100 engines operating under a single operating condition with varying fault progressions. To facilitate anomaly detection, we implement a 60/40 split, as suggested by similar studies in the literature [18, 19], where the first 60% of cycles for each engine are labeled as normal, and the remaining 40% are considered anomalous. Following this segmentation, we apply a random 80/20 train-test split across all engines to ensure a robust evaluation of our anomaly detection approach.
@@
The dataset includes engine operational settings, temperature measurements at various engine components, pressure readings, fan and core speed indicators, as well as different ratios and flow parameters related to efficiency and fuel usage. Additionally, variables associated with coolant bleed and bleed enthalpy provide insights into engine cooling and overall performance.
III-C2 Applying the Theoretical Foundation
@@
In the NASA C-MAPSS dataset, the measurement space has a dimension of $\mathbf{k=24}$, representing sensor readings from the turbine system. To determine a suitable embedding dimension consistent with Th. 1, we make the following assumptions about the system’s essential dynamics.
We hypothesize that the turbine’s core behavior is captured by three one-dimensional attractors, which could correspond to three state-derivative pairs, such as:
  • Rotational speed and its rate of change,
  • Temperature and its rate of change,
  • Pressure and its rate of change.
Given that the measurement function provides a structured representation of the underlying physical dynamics, we assume that it maps the real attractor dimension 3 to a box-counting dimension below 4, Based on Th. 1, this allows us to select an embedding dimension of $\mathbf{n=8}$, ensuring a one-to-one mapping.
@@
Based on our approach, we select four state-derivative pairs as the embedding. To ensure that this representation is at least an immersion in the subspace corresponding to the system’s normal or stable regime, we apply a numerical method to verify whether the Jacobian has full rank in the normal state of the test data. While this does not constitute a strict mathematical proof, it serves as a reliable guideline for assessing the validity of the embedding.
Singular Value Decomposition (SVD) is a fundamental matrix factorization technique that can be applied to any complex-valued matrix $X\in\mathbb{C}^{n\times m}$. The decomposition expresses $X$ in the following form:
$X=U\Sigma V^{*},$
where:
  • $U\in\mathbb{C}^{n\times n}$ is a unitary matrix with orthonormal columns.
  • $\Sigma\in\mathbb{R}^{n\times m}$ is a diagonal matrix with nonnegative real numbers on the diagonal and zeros elsewhere. These diagonal entries are known as the singular values of $X$.
  • $V^{*}\in\mathbb{C}^{m\times m}$ is the conjugate transpose of a unitary matrix $V\in\mathbb{C}^{m\times m}$.
This decomposition exists for any complex matrix and provides significant insights into the structure of the matrix. The singular values in $\Sigma$ are ordered such that:
$\sigma_{1}\geq\sigma_{2}\geq\dots\geq\sigma_{\min(n,m)}\geq 0.$
SVD is particularly powerful for analyzing the rank of a matrix. The rank of $X$ can be determined by counting the number of non-zero singular values in $\Sigma$. If all singular values are non-zero, the matrix is full rank [28].
III-C3 Temporal Differential Consistency informed Autoencoder

### v2

@@ Abstract
Anomaly detection in complex dynamical systems is essential for ensuring reliability, safety, and efficiency in industrial and cyber-physical infrastructures. Predictive maintenance helps prevent costly failures, while cybersecurity monitoring has become critical as digitized systems face growing threats. Many of these systems exhibit oscillatory behaviors and bounded motion, requiring anomaly detection methods that capture structured temporal dependencies while adhering to physical consistency principles. In this work, we propose a system-theoretic approach to anomaly detection, grounded in classical embedding theory and physics-inspired consistency principles. We build upon the Fractal Whitney Embedding Prevalence Theorem that extends traditional embedding techniques to complex system dynamics. Additionally, we introduce state-derivative pairs as an embedding strategy to capture system evolution. To enforce temporal coherence, we develop a Temporal Differential Consistency Autoencoder (TDC-AE), incorporating a TDC-Loss that aligns the approximated derivatives of latent variables with their dynamic representations. We evaluate our method on the C-MAPSS dataset, a benchmark for turbofan aeroengine degradation. TDC-AE outperforms LSTMs and Transformers while achieving a nearly 200x reduction in MAC operations, making it particularly suited for lightweight edge computing. Our findings support the hypothesis that anomalies disrupt stable system dynamics, providing a robust signal for anomaly detection.
@@ I Introduction
Anomaly detection in complex physical dynamical systems is a critical research area as industrial and engineered systems become more sophisticated. Identifying deviations from expected behavior is essential for ensuring reliability, safety, and efficiency. This is particularly relevant in predictive maintenance and cybersecurity monitoring. In industrial systems and rotating machinery, early fault detection helps prevent costly failures and downtime [1]. Meanwhile, as critical infrastructures, such as power grids and water distribution systems, become increasingly digitized, the risk of cyber threats has grown significantly [2, 3]. Recent attacks [4, 5] on cyber-physical systems highlight the need for robust monitoring techniques to detect both malicious intrusions and system failures. Ensuring the security and stability of these dynamical systems requires adaptive anomaly detection methods capable of addressing evolving threats and operational challenges.
@@ II-A Time-Series Modeling
Models like LSTMs and Transformers with attention mechanisms have been successfully applied to time-series anomaly detection in fields such as mechanical engineered system, aerospace, and industrial monitoring [6, 7, 8]. While these methods effectively capture long-range dependencies and irregular temporal patterns, they are computationally expensive. For example, evaluated on common benchmark datasets, sequence lengths up to 500 time steps were required to achieve competitive performance [9]. The high memory requirements often exceed the constraints of typical MCU-level devices, making real-time deployment challenging. For example, a typical LSTM with 500 time steps and 50 hidden units already surpasses the available memory of common low-power devices, limiting its practical applicability. Additionally, the sequential nature of RNNs restricts parallelization, further increasing computational cost [10]. These challenges highlight the need for alternative approaches that balance computational efficiency with robust anomaly detection, aligning with broader goals of sustainability and practical deployability.
@@ II-B Physics-Informed Methods
In many applications, incorporating domain knowledge can help reduce computational demands. Physics-informed neural networks (PINNs) embed physical laws directly into neural networks, enabling them to leverage known system dynamics [11, 12, 13]. However, applying PINNs to complex dynamical physical systems remains challenging. These systems involve numerous interdependent physical principles, and explicitly modeling them within a neural network would require domain expertise, making large-scale applications impractical [14, 15].
@@ II-D Proposed Approach
Instead of relying on complex models to differentiate normal and abnormal behavior, we adopt a system-theoretic approach inspired by classical embedding theory to analyze the system’s latent dimensionality. By incorporating physics-inspired consistency principles, we approximate the underlying causal mechanisms without explicitly enforcing physical laws. We hypothesize that in a stable regime, complex systems exhibit predictable behavior, enabling well-approximated low-dimensional embeddings. In contrast, anomalies introduce additional complexity, disrupting these stable relationships. We aim to leverage this property to detect abnormal states when the learned embedding no longer adequately captures the system dynamics.
@@ III Theoretical Foundation & Methods
We study the mathematical foundations of embedding bounded dynamical systems into an autoencoder’s latent space, aiming to ensure reliable mapping of system dynamics for anomaly detection and health monitoring.
Consider for example an oscillatory dynamical system, specifically a damped-driven pendulum. Initially, the system remains within a bounded oscillatory regime; however, due to the influence of drift terms, it gradually deviates, eventually departing from the confined state space. This perturbation effect is illustrated in Figure 1, where the phase space trajectory initially remains within a bounded region but progressively transitions into an unbounded state. This formulation highlights that the state variables of an oscillatory system, namely the angular displacement $\theta$ and its derivative $\dot{\theta}$, define a phase space where the system’s behavior can be systematically analyzed. Within this phase space, bounded regions correspond to normal oscillations, where the system follows a predictable, periodic trajectory. However, once the system leaves this bounded region, it signifies the onset of unbounded oscillations or deviations from regular behavior. For formal definitions and details on the system’s implementation, refer to Appendix A.
@@
We explore whether this concept extends to complex dynamical systems and the constraints and conditions for a successful transition. The challenge lies in preserving phase space properties in high-dimensional, nonlinear systems while maintaining topological consistency and dynamical coherence. To address this, we examine the mathematical and computational requirements for embedding system dynamics into a reduced representation, guiding the design of effective autoencoder-based anomaly detection.
III-A Embedding Complex System Dynamics into a Latent Space
We begin by formalizing the mathematical foundation for embedding high-dimensional dynamical systems into a reduced representation. Consider a complex bounded dynamical system with state variables $x\in\mathbb{R}^{m}$, where $m$ represents the dimension of the physical system and is typically much larger than the number of available measurements $(m\gg k$).
@@ III-A Embedding Complex System Dynamics into a Latent Space
We analyze the latent space encoding of a dynamical system from a systems-theoretic perspective. Ideally, such systems in a stable state evolve on smooth manifolds, but real-world dynamics often unfold on complex, non-smooth sets due to measurement noise and the inherent need for machines to exhibit more intricate dynamics to perform their functions effectively. A key aspect of this generalization is the notion of dimensionality for compact subsets, which leads to the definition of box-counting dimension.
@@
Let $A$ be a bounded subset of $\mathbb{R}^{k}$. For each $\epsilon>0$, let $N_{\epsilon}(A)$ denote the smallest number of sets of diameter at most $\epsilon$ required to cover $A$.
The box-counting dimension of $A$ is defined as
$\dim_{B}(A)=\lim_{\epsilon\to 0}\frac{\ln N_{\epsilon}(A)}{\ln(1/\epsilon)}$
provided the limit exists.
In this context, $\epsilon$ provides a measure of the complexity exhibited by motion beyond smooth manifold assumptions.
@@
We discuss the necessary assumptions to apply this formalism to our system representation.
Compactness of the System State: We assume that the key drivers of the physical process form a bounded and closed subset of $\mathbb{R}^{k}$, ensuring compactness. This assumption reflects the practical reality that real-world physical systems with inherently bounded dynamics, when operating in a stable state, remain confined within a finite region of their state space.
Informative Measurement Function: We assume that the measurement function retains sufficient information about the system dynamics. Specifically, it must be injective enough over $A$ to avoid collapsing distinct states, ensuring that the measurement space captures the necessary complexity of the system.
Dimensionality of the System: We assume that the measurement-level set $A\subset\mathbb{R}^{m}$ has a well-defined box-counting dimension $d$, which reflects the effective complexity of the system’s dynamics. This dimensionality determines the necessary embedding space.
Under these conditions, the theorem is applicable. This allows us to estimate the required embedding dimension $n$ to achieve a faithful description of the system’s dynamics.
@@
To effectively utilize this type of embedding, we need more than just a one-to-one mapping—we also require consistent differential structure. In other words, small changes in the latent space should produce meaningful, non-degenerate changes in the measurement space. A full-rank Jacobian is a necessary and sufficient condition to guarantee this kind of well-behaved, locally reversible mapping [23].
@@
III-B Application to Autoencoder-based Representations
With the theoretical framework established, we apply it to a practical case in two steps. First, we select a stable, bounded dynamical system with accessible measurements. For this, we use the NASA C-MAPSS dataset, which provides time-series sensor data from aircraft engines. Second, we develop an autoencoder that learns embeddings aligned with our framework, ensuring a one-to-one mapping and immersion within the normal operating regime of the system.
III-B1 C-MAPSS Dataset
The NASA Commercial Modular Aero-Propulsion System Simulation (C-MAPSS) dataset [24] is a widely used benchmark in prognostics and health management studies [25, 26, 27].
The dataset consists of multivariate time series capturing engine operation under varying initial wear, manufacturing differences, and noisy sensor readings. Each engine starts in a normal state before gradually developing a fault.
@@ III-B1 C-MAPSS Dataset
We use the FD001 subset of the C-MAPSS dataset, which contains run-to-failure time series from 100 independent engines operating under a single operating condition but exhibiting varying fault progressions. This substantial number of individual systems allows us to assess the generalization capability of our method across multiple degradation trajectories. For anomaly detection, we follow a 60/40 split, where the first 60% of cycles are labeled as normal and the remaining 40% as anomalous, as suggested in similar studies [18, 19]. A random 80/20 train-test split based on separate engines ensures robust evaluation.
@@
The dataset includes sensor readings for temperature, pressure, fan and core speed, efficiency ratios, fuel flow, and coolant bleed parameters, providing insights into engine performance and degradation.
III-B2 Applying the Theoretical Foundation
@@
The NASA C-MAPSS dataset provides a measurement space of dimension $k=24$, representing sensor readings from the turbine system. To determine a suitable embedding dimension in accordance with Theorem 1, we establish the following assumptions about the system’s essential dynamics.
As argued in the theory section, the turbine operates in a stable regime, where its state evolution remains within a well-defined, physically constrained region. This ensures that the system’s key drivers form a bounded and closed subset of $\mathbb{R}^{k}$, thereby satisfying the compactness assumption.
We hypothesize that the turbine’s core behavior is governed by three fundamental state-derivative pairs, corresponding to rotation, temperature, and pressure along with their respective rates of change. These represent the system’s intrinsic dynamics and serve as a basis for estimating its fractal dimension. In Appendix B, we computed the box-counting dimension numerically for the attractor of a damped driven pendulum, yielding $d\approx 1.3$. For a perfect mathematical pendulum, this value reduces to $d=1$, reflecting its simple periodic structure (a circle). Given the structural similarities between the fundamental dynamics of the pendulum and the turbine’s primary state variables, we assume a similar level of dynamic complexity here. However, this framework provides room to incorporate domain knowledge, such as machine tolerances and operational constraints, to refine this assumption further. This leads to an estimated box-counting dimension $d<4$ for the combined system. Applying Theorem 1, this result suggests the selection of an embedding dimension of $n=8$.
The measurement function in the C-MAPSS dataset is assumed to retain sufficient information about the system’s essential dynamics. Given the structured nature of the sensor readings, it provides a reliable representation of the turbine’s state evolution without collapsing distinct states, thereby satisfying the injectivity condition.
@@
Based on our approach, we select four state-derivative pairs for the embedding. To ensure that this representation is at an immersion in the subspace corresponding to the system’s normal or stable regime, we apply a numerical method to verify whether the Jacobian has full rank in the normal state of the test data. While this does not constitute a strict mathematical proof, it serves as a reliable guideline for assessing the validity of the embedding. SVD is particularly powerful for analyzing the rank of a matrix. The rank of $X$ can be determined by counting the number of non-zero singular values in $\Sigma$. If all singular values are non-zero, the matrix is full rank [28].
Theorem 1 assumes a smooth map and that it applies to ”almost every” map. A $\tanh$-activated neural network encoder reasonably satisfies the smoothness assumption. The ”almost every” condition applies to all smooth maps, whereas neural networks form a restricted subset. However, by universal approximation, a sufficiently large network can approximate such an embedding, making the theorem a strong and relevant practical guideline.
III-B3 Temporal Differential Consistency informed Autoencoder